# On the “verification” of a Nicolas inequality

### v5

@@ Abstract.
Note to the reader
First version of the present preprint has an error that brings us to an incorrect conclusion. In second version we amend this error, but we do not provide new contributions. In this new version we have corrected Step 1 of final discussion and added Remark 2.4, and also (as previously) we would advise the reader that the paper contains some imprecisions, that the author have left to the reader, and also could include some undetected mistakes, so the author appreciates any comment without attached files (in order to avoid fishing) to galdames@uv.es .
@@ 1. Introduction and notation
Nicolas inequality was introduced in [6, Theorem 2] and establish that one and only one of the following holds:
@@ 2. On the sign of the error term
Taking into account Mertens Theorem (see e.g. [7, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
@@
Observe that $P_{x}(n)$ is increasing in $x$ and converges to $P(n)$, thus for each $\varepsilon>0$ there exists $x_{1}$ such that $P(n)-P_{x}(n)<\varepsilon$ for every $x\geq x_{1}$. In addition one can deduce that
@@
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant ([9, (2.7)]).
@@
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [9, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
@@
Now we take into account computations of (2), equation (3), comments above it and add the Euler constant $\gamma$ to both sides of this inequality for $x\geq x_{3}=\max\{x_{0},x_{1},x_{2}\}$
@@
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [6, Theorem 2] there are only two possibilities, (T) must hold.
@@ 2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [7, Lemma 8.2 and Theorem 8.8(a)] and also [2]).
@@
Step 2. We follow with a sum involving the Mangoldt function (see [7, Theorem 8.8(a)]).
@@
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Taking into account [7, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ for a sufficient large integer $x$, thus
$\begin{split}\sum_{d\leq x}\Big\{\frac{x}{d}\Big\}\Lambda(d)&=\sum_{d\leq x/2}\Big\{\frac{x}{d}\Big\}\Lambda(d)+\sum_{x/2<d<x}\frac{1}{d}\Lambda(d)\\ &<\sum_{d\leq x/2}\Lambda(d)+\frac{2}{x+1}\sum_{x/2<d<x}\Lambda(d)\\ &<\psi(x/2)+\frac{2}{x+1}(\psi(x)-\psi(x/2))\\ &<\frac{3}{4}\psi(x)+\frac{2}{x+1}\psi(x)=\frac{3x+11}{4x+4}\psi(x).\end{split}$
Moreover, if we avoid the terms for $d\leq x/2$ in the first equality, we have
$\sum_{d\leq x}\Big\{\frac{x}{d}\Big\}\Lambda(d)>\frac{1}{x}\big(\psi(x)-\psi(x/2)\big)\,.$
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=1.170533$..., for a sufficiently large $x$ (see [7, Theorem 8.3]), we deduce that
$\frac{0.1950884}{x}<\frac{\displaystyle\sum_{d\leq x}\Big\{\dfrac{x}{d}\Big\}\Lambda(d)}{x}<0.58526625\,$,
where 0.1950884<ac/20.1950884<a-c/2 and $0.58526625<\dfrac{3}{4}a$. On the other hand, as $O_{1}(\log x)/x$ is negative and converges to zero, there is a $x_{2}$ such that $0<O_{2}(1)<0.585$ for every $x\geq x_{2}$.
Step 3. Next we deal with a sum over primes (see [7, Theorem 8.8(b)]).
@@
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [7, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
@@
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [7, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
@@
Remark $2.$
Observe that from approximations (4) and (5) we have that
$|O(1/\log x)|<0.02/\log x\,$,
for sufficiently large $x$.
Remark 2.3.
@@
$\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=-b-\log\log 2+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,.$
However, from these equations we just can deduce that $\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}=\log\log 2-\log\log x$.
Remark 2.4.
Assume that
$O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=\frac{C(x)}{\log x}\,$,
thus
$\begin{split}C(x)&=\log x\cdot O(1/\log x)=R(x)-\log x\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\\ &=\sum_{p\leq x}\frac{\log p}{p}-\log x\Big(1+\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\Big)\end{split}$
and $\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt$ goes to zero when $x$ grows to infinity. In consequence $C(x)$ goes to $R(x)$ that is negative and bounded from above by a negative number (see (6)), that is $C(x)<C$ for some $C<0$

### v6

# On the verification of a Nicolas inequality

@@ Abstract.
Author’s note
The sequence of versions of this paper that we have reviewed illustrates the evolution from an astonishing result to another one that could be a good tool for studying Nicolas inequality. We have been able to trace the errors, mistakes, typos and inaccuracies from a first draft to a final version. Despite this does not guarantee the paper is entirely free of oversights.
@@ 1. Introduction and notation
Nicolas inequality was introduced in [7, Theorem 2] and establish that one and only one of the following holds:
@@ 2. On the sign of the error term
Taking into account Mertens Theorem (see e.g. [8, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
@@
Observe that $P_{x}(n)$ is increasing in $x$ and converges to $P(n)$, thus for each $\varepsilon>0$ there exists $x_{1}$ such that
(3) $\sum_{n=2}^{\infty}\dfrac{P(n)}{n}-\sum_{n=2}^{\infty}\dfrac{P_{x}(n)}{n}<\varepsilon\,,\qquad\mbox{ for every }x\geq x_{1}\,.$
In addition one can deduce that
@@
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant ([10, (2.7)]).
@@
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [10, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
@@
Now we take into account computations of (2), equation (4), (3) and add the Euler constant $\gamma$ to both sides of this inequality for $x\geq x_{3}=\max\{x_{0},x_{1},x_{2}\}$
@@
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [7, Theorem 2] there are only two possibilities, (T) must hold.
@@ 2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [8, Lemma 8.2 and Theorem 8.8(a)] and also [2]).
@@
Step 2. We follow with a sum involving the Mangoldt function (see [8, Theorem 8.8(a)]).
@@
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Analogously for $d\in(x/3,x/2)$, the reminder is $2$, so in this case $\Big\{\dfrac{x}{d}\Big\}=\dfrac{2}{d}$ and again $\Big\{\dfrac{x/2}{x/2}\Big\}=0$. Taking into account [8, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ and $\psi(x/3)<1/2\psi(x)$ for a sufficient large integer $x$, thus
$\begin{split}\sum_{d\leq x}\Big\{\frac{x}{d}\Big\}\Lambda(d)&=\sum_{d\leq x/3}\Big\{\frac{x}{d}\Big\}\Lambda(d)+\sum_{x/3<d\leq x/2}\frac{2}{d}\Lambda(d)+\sum_{x/2<d\leq x}\frac{1}{d}\Lambda(d)\\ &<\sum_{d\leq x/2}\Lambda(d)+\frac{6}{x+1}\sum_{x/3<d<x/2}\Lambda(d)+\frac{2}{x+1}\sum_{x/2<d<x}\Lambda(d)\\ &<\psi(x/3)+\frac{6}{x+1}(\psi(x/2)-\psi(x/3))+\frac{2}{x+1}(\psi(x)-\psi(x/2))\\ &<\frac{1}{2}\psi(x)+\frac{6}{x+1}\psi(x)+\frac{2}{x+1}\psi(x)=\frac{x+17}{2x+2}\psi(x).\end{split}$
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=3/2a$, for a sufficiently large $x$ (see [8, Theorem 8.3]), we deduce that
$0<\frac{\displaystyle\sum_{d\leq x}\Big\{\dfrac{x}{d}\Big\}\Lambda(d)}{x}<0.58526625\,$,
where $0.58526625<\dfrac{3}{4}a$. On the other hand, as $O_{1}(\log x)/x$ converges to zero, there is a $x_{2}$ such that $1<O_{2}(1)<-0.415=0.585-1$ for every $x\geq x_{2}$.
Step 3. Next we deal with a sum over primes (see [8, Theorem 8.8(b)]).
@@
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [8, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
@@
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [8, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
@@
$O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=-b-\log\log 2+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,.$
Conclusion
Observe that $O(1/\log x)$ converges to zero and, from the last equation, it is the difference between an step function with discontinuities on primes and an strictly monotone function. We already saw this kind of functions in [9] (for instance), namely $\pi(x)-\operatorname{li}(x)$, $\psi(x)-x$ and $\theta(x)-x$, and we know they change of sign infinitely asymptoticaly. So it seems reasonable to think that $O(1/\log x)$ also change of sign intinitely for sufficiently large $x$, but this is another issue we do not deal here (see [4, Chapter V] for more information on these irregularities).
Acknowledgment
We would like to thank Professors Wladyslaw Narkiewicz and Enrique A. Sánchez Pérez for their comments and suggestions.